Question 3: If the domain of 𝑓(π‘₯)=3π‘Žπ‘₯+22𝑏π‘₯βˆ’1 is given by 𝐷={π‘₯|π‘₯βˆˆβ„,π‘₯β‰ 3} and the graph of this function passes through the point (βˆ’3,8). Write down the values of π‘Ž and 𝑏.


1
Expert's answer
2021-04-06T16:59:15-0400

For a given function



"f(x)=\frac{3ax+2}{2bx-1}\to\,\,\,D=\left\{x\left|2bx-1\neq0\right.\right\}\\[0.3cm]
D=\left\{x\left|x\neq\frac{1}{2b}\right.\right\}\,\,\,\text{and}\,\,\,D=\left\{x\left|x\in\mathbb{R},x\neq3\right.\right\}\left(\text{by condition}\right)\\[0.3cm]
\frac{1}{2b}=3\to\boxed{b=\frac{1}{6}}"



By the condition of the problem, the function "f(x)" passes through the point "(-3,8)", which means that



"f(-3)=8\to\frac{2a\cdot(-3)+2}{2\cdot\displaystyle\frac{1}{6}\cdot(-3)-1}=8\to\\[0.3cm]
\frac{2\cdot(-3a+1)}{-2}=8\to3a-1=8\to a=\frac{9}{3}=3\to\boxed{a=3}"

Conclusion,



"\boxed{a=3\,\,\,\text{and}\,\,\,b=\frac{1}{6}}"

ANSWER



"a=3\,\,\,\text{and}\,\,\,b=\frac{1}{6}"

Q.E.D.